### v2

@@ 2.2 Competitive equilibrium
In a Fisher market instance, given is a set of agents ${\mathcal{A}}$ and a set of divisible goods ${\mathcal{M}}$, with one unit available from each good. Agents have utility functions $u_{i}\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ and budgets $b_{i}\in\mathbb{R}_{>0}$, $i\in{\mathcal{A}}$; we denote the budget vector as $\mathbf{b}=(b_{i})_{i\in{\mathcal{A}}}$. An allocation vector is $\mathbf{x}\in\mathbb{R}_{\geq 0}^{{\mathcal{A}}\times{\mathcal{M}}}$, where $x_{ij}$ denotes the amount of good $j$ allocated to agent $i$. A price vector is a vector $\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$. The demand correspondence of agent $i$ will be denoted as ${\mathcal{D}_{i}}(\mathbf{p},b_{i})$.
@@
Given a Fisher market instance as above, the allocations and prices $(\mathbf{x},\mathbf{p})$ form a competitive (market) equilibrium if the following hold:
@@
    every agent gets an optimal utility at these prices: $x_{i}\in{\mathcal{D}_{i}}(\mathbf{p},b_{i})$ for every agent $i\in{\mathcal{A}}$
@@
    every item with positive price is fully sold: $\sum_{i}x_{ij}=1$ if $p_{j}>0$.
@@
Given a Fisher market instance and $\alpha\geq 1$, the allocations and prices $(\mathbf{x},\mathbf{p})$ form a $\alpha$-demand approximate competitive equilibrium if
@@
    $\left\langle\mathbf{p},x_{i}\right\rangle\leq b_{i}$ for each $i\in{\mathcal{A}}$, and $i$ gets at least $\alpha$-fraction of the optimal utility at these prices: $u_{i}(x_{i})\geq\frac{1}{\alpha}\max_{x_{i}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\{u_{i}(x_{i})\,:\,\left\langle\mathbf{p},x_{i}\right\rangle\leq b_{i}\}$.
@@
    every item with positive price is fully sold: $\sum_{i}x_{ij}=1$ if $p_{j}>0$.
It is a well-established fact that when all $b_{i}$’s are identical, the competitive equilibrium attains exact envy-freeness. Here, we demonstrate that our definition of approximate competitive equilibrium also results in approximate envy-freeness.
@@
For a given allocation $\mathbf{x}\in\mathbb{R}_{\geq 0}^{{\mathcal{A}}\times{\mathcal{M}}}$, we say it is $\beta$-envy-free if, for any agents $i$ and $j$, the condition $u_{i}(x_{i})\geq\frac{1}{\beta}u_{i}(x_{j})$ holds.
@@
When $b_{i}$’s are identical, the allocation $\mathbf{x}$ of an $\alpha$-approximate competitive equilibrium is also $\alpha$-envy-free.
@@
This follows directly from $u_{i}(x_{i})\geq\frac{1}{\alpha}\max_{x_{i}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\left\{u_{i}(x_{i})\,:\,\left\langle\mathbf{p},x_{i}\right\rangle\leq b_{i}\right\}\geq\frac{1}{\alpha}u_{i}(x_{j})$. ∎
@@
In case there are satiable utility functions, $\left\langle\mathbf{p},x_{i}\right\rangle<b_{i}$ may be possible in an (exact) competitive equilibrium. Hylland and Zeckhauser (1979) showed that competitive equilibrium may not be Pareto efficient due to satiable utilities. Mas-Colell (1992) strengthened the definition of a competitive equilibrium to require that each agent choose a least-cost optimal plan. This is also known as a strong equilibrium.
@@
Given a Fisher market instance, the allocations and prices $(\mathbf{x},\mathbf{p})$ form a thrifty competitive equilibrium if $(\mathbf{x},\mathbf{p})$ is a competitive equilibrium, and moreover, for every $i\in{\mathcal{A}}$, the allocation $x_{i}$ minimizes $\left\langle\mathbf{p},x_{i}\right\rangle$ among all allocations ${\mathcal{D}_{i}}(\mathbf{p},b_{i})$.
@@
Note that in case all utilities are non-satiable, competitive equilibria and thrifty competitive equilibria coincide.
@@ 2.3 Nash welfare
Given a Fisher market instance, the Nash welfare of an allocation $\mathbf{y}\in\mathbb{R}_{\geq 0}^{{\mathcal{A}}\times{\mathcal{M}}}$ is the weighted geometric mean of their utilities, namely,
$\left(\prod_{i\in{\mathcal{A}}}u_{i}(x_{i})^{b_{i}}\right)^{\frac{1}{\sum_{i\in{\mathcal{A}}}b_{i}}}\,.$
The allocation $\mathbf{y}\in\mathbb{R}_{\geq 0}^{{\mathcal{A}}\times{\mathcal{M}}}$ maximizes the (weighted) Nash welfare, if its Nash welfare is maximal along all allocations with $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1$ for all $j\in{\mathcal{M}}$, $\mathbf{y}\geq\mathbf{0}$.
Note that the Nash welfare maximizing allocations are precisely the optimal solutions to Eisenberg–Gale convex program Eisenberg and Gale (1959).
[…]
@@
$\varphi(\mathbf{q})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{i\in{\mathcal{A}}}\max_{y_{i}\in\mathbb{R}_{\geq 0}^{{\mathcal{M}}}}\left(b_{i}\log u_{i}(y_{i})-\left\langle\mathbf{q},y_{i}\right\rangle\right)+\sum_{j\in{\mathcal{M}}}q_{j}.$ (4)
We define the domain $\operatorname{dom}(\varphi)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\cap_{i\in{\mathcal{A}}}\operatorname{pdom}(u_{i})$ (see (3)). Thus, $\operatorname{dom}(\varphi)$ is the set of nonnegative prices $\mathbf{q}$ such that $\varphi(\mathbf{q})<\infty$. Strong duality holds and the dual optimum is taken according since Slater’s condition holds. Consequently, the optimal Nash welfare value equals the dual optimum
$\min_{\mathbf{q}\in\operatorname{dom}(\varphi)}\varphi(\mathbf{q})\,.$ (5)
This optimum is taken (see Lemma 6.2), and we refer to the optimal $\mathbf{q}\in\operatorname{dom}(\varphi)$ as Gale prices. We note that these may not be unique. To justify this term, let us recall the definition of the Gale demand correspondence (2). The term corresponding to agent $i$ in the definition of $\varphi(\mathbf{q})$ is the value of the assignments in the Gale demand correspondence of this agent at prices $\mathbf{q}$ and budget $b_{i}$. We use the notation $\mathcal{GD}_{i}(\mathbf{q},b_{i})$ for the Gale demand for the utility $u_{i}$, $i\in{\mathcal{A}}$.
@@ 2.4 Gale substitute utilities
    for any $b>0$, any $\mathbf{q},\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$, and any $y\in\mathcal{GD}^{u}(\mathbf{q},b)$, if $\mathbf{q}\geq\mathbf{q}^{\prime}$, then there exists a $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ such that $y^{\prime}_{j}\leq y_{j}$ whenever $q_{j}^{\prime}=q_{j}$;
@@
    for any $b\geq b^{\prime}>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$, and for any $y\in\mathcal{GD}^{u}(\mathbf{q},b)$, there exists a $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$ such that $y^{\prime}\leq y$.
We say that $u(\cdot)$ is $\Sigma$-Gale-substitutes, if it is Gale-substitutes and the following additional property holds:
$\mbox{if }\mathbf{0}\in\operatorname{pdom}(u)\,,\,b>0\mbox{ and }\mathbf{q}\in\mathbb{R}_{\geq 0}^{\mathcal{M}},\mbox{ then for any }y\in\mathcal{GD}(\mathbf{0},b),\exists y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b)\,:\,y^{\prime}\leq y\,.$ ($\Sigma$)
@@
As noted above, $\mathbf{0}\in\operatorname{pdom}(u)$ if and only if $u$ is satiable; thus, ($\Sigma$) is vacuously true for non-satiable utilities. In the satiable case, it asserts that if $y$ is a bundle of maximum possible utility, then at any prices there is an optimal bundle below $y$. Observe that this is immediate for separable utilities. On the other hand, ($\Sigma$) fails for the simple Leontief-free utility $u(x)=\max\{x_{1}+x_{2},1\}$: $y=(1,0)$ is an optimal bundle, whereas for $b=1$, $\mathbf{q}=(2,1)$, the unique solution in $\mathcal{GD}^{u}(\mathbf{q},b)$ is $y^{\prime}=(0,1)$.
Next, we demonstrate a utility function that exhibits Gale-substitutes but not substitutes properties.
@@
Consider an agent with budget $b$ and a utility function over two goods given by $u(x)=\min\{x_{1},0.4\}+x_{2}$. This utility is separable and, as such, Gale-substitutes according to Lemma 7.1 shown later. However, we show that this is not weak gross substitutes (WGS): there exist prices $\mathbf{p}\leq\mathbf{p}^{\prime}$ and $x\in{\mathcal{D}}^{u}(\mathbf{p},b)$ such that for some good $j$ with $p_{j}=p^{\prime}_{j}$, every $x^{\prime}\in{\mathcal{D}}^{u}(\mathbf{p},b)$ has $x^{\prime}_{j}<x_{j}$. That is, let $\mathbf{p}=(0.8,1)$ and $\mathbf{p}^{\prime}=(0.9,1)$, and $b=1$. It is easy to see that $x=(0.4,0.68)$ is the unique optimal bundle at prices $\mathbf{p}$, and $x^{\prime}=(0.4,0.64)$ is the unique optimal bundle at prices $\mathbf{p}^{\prime}$ in contradiction to WGS property.
3 Overview of techniques
The ‘price of anarchy’ result Theorem 1.1 is proved using an elementary argument on concavity and the log-sum inequality. The proof is inspired by an argument in Theorem 1 in Brânzei et al. (2022). Example A.4 shows that the bound $(1/\mathrm{e})^{1/\mathrm{e}}$ is tight.
To prove Theorem 1.2 showing that the Nash welfare maximizing allocation is 2-demand-approximate CE, the key ingredient is Lemma 5.1, showing that the solutions in the Gale demand system are within budget, that is, for $y\in\mathcal{GD}^{u}(\mathbf{q},b)$ we have $\left\langle\mathbf{q},y\right\rangle\leq b$. Since at the Nash welfare maximizing solution, every agent gets an allocation from the Gale demand system at the dual prices $\mathbf{q}$, this lemma guarantees that the allocation is feasible. The 2-demand-approximate property then follows by a supergradient argument.
The proof of Theorem 1.3 relies on two key lemmas derived using Lagrangian duality and supergradient arguments. The first is a simple observation, showing that the Gale demand solution for lower prices yields at least half the utility of the utility maximizing bundle at the current prices.

### v3

@@ 2.2 Competitive equilibrium
In a Fisher market instance, we are given a set of agents $\mathcal{A}$} and a set of divisible goods $\mathcal{M}$}, with one unit available of each good. Agents have utility functions $u_{i}\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ and budgets $b_{i}\in\mathbb{R}_{>0}$, $i\in{\mathcal{A}}$; we denote the budget vector as $\mathbf{b}=(b_{i})_{i\in{\mathcal{A}}}$. An allocation is represented by $(\mathbf{x}_{i})_{i\in{\mathcal{A}}}$, where $x_{ij}$ denotes the amount of good $j$ allocated to agent $i$. A price vector is a vector $\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$. The demand correspondence of agent $i$ will be denoted as ${\mathcal{D}_{i}}(\mathbf{p},b_{i})$.
@@
Given a Fisher market instance as above, the allocations and prices $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$ form a competitive (market) equilibrium if the following hold:
@@
    every agent gets an optimal utility at these prices: $\mathbf{x}_{i}\in{\mathcal{D}_{i}}(\mathbf{p},b_{i})$ for every agent $i\in{\mathcal{A}}$.
@@
    every good with positive price is fully sold: $\sum_{i}x_{ij}=1$ if $p_{j}>0$.
@@
Given a Fisher market instance and $\alpha\geq 1$, the allocations and prices $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$ form an $\alpha$-demand approximate competitive equilibrium if
@@
    $\left\langle\mathbf{p},\mathbf{x}_{i}\right\rangle\leq b_{i}$ for each $i\in{\mathcal{A}}$, and $i$ gets at least $\alpha$-fraction of the optimal utility at these prices: $u_{i}(\mathbf{x}_{i})\geq\frac{1}{\alpha}\max_{\mathbf{x}_{i}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\{u_{i}(\mathbf{x}_{i})\,:\,\left\langle\mathbf{p},\mathbf{x}_{i}\right\rangle\leq b_{i}\}$.
@@
    every good with positive price is fully sold: $\sum_{i}x_{ij}=1$ if $p_{j}>0$.
It is well known that when all $b_{i}$’s are identical, competitive equilibria yield an allocation that is exactly envy-free. Here, we demonstrate that our definition of approximate competitive equilibrium also results in approximate envy-freeness.
@@
For a given allocation $(\mathbf{x}_{i})_{i\in{\mathcal{A}}}$, we say it is $\beta$-envy-free if, for any agents $i$ and $j$, the condition $u_{i}(\mathbf{x}_{i})\geq\frac{1}{\beta}u_{i}(\mathbf{x}_{j})$ holds.
@@
When $b_{i}$’s are identical, the allocation $(\mathbf{x}_{i})_{i\in{\mathcal{A}}}$ of an $\alpha$-approximate competitive equilibrium is also $\alpha$-envy-free.
@@
This follows directly from $u_{i}(\mathbf{x}_{i})\geq\frac{1}{\alpha}\max_{\mathbf{x}_{i}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\left\{u_{i}(\mathbf{x}_{i})\,:\,\left\langle\mathbf{p},\mathbf{x}_{i}\right\rangle\leq b_{i}\right\}\geq\frac{1}{\alpha}u_{i}(\mathbf{x}_{j})$. ∎
@@
In the case of satiable utility functions, it is possible for an (exact) equilibrium to satisfy $\left\langle\mathbf{p},\mathbf{x}_{i}\right\rangle<b_{i}$. Hylland and Zeckhauser (1979) showed that competitive equilibria may not be Pareto efficient in the presence of satiable utilities. Mas-Colell (1992) strengthened the definition of competitive equilibrium by requiring that each agent select a least-cost optimal bundle. This is also known as a strong equilibrium.
@@
Given a Fisher market instance, the allocations and prices $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$ form a thrifty competitive equilibrium if $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$ is a competitive equilibrium, and moreover, for every $i\in{\mathcal{A}}$, the allocation $\mathbf{x}_{i}$ minimizes $\left\langle\mathbf{p},\mathbf{x}_{i}\right\rangle$ among all allocations ${\mathcal{D}_{i}}(\mathbf{p},b_{i})$.
@@
Note that when all utility functions are non-satiable, every competitive equilibrium is thrifty.
@@ 2.3 Nash welfare
Given a Fisher market instance, the Nash welfare of an allocation $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$ is the weighted geometric mean of their utilities, namely,
$\left(\prod_{i\in{\mathcal{A}}}u_{i}(\mathbf{y}_{i})^{b_{i}}\right)^{\frac{1}{\sum_{i\in{\mathcal{A}}}b_{i}}}\,.$
The allocation $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$ maximizes the (weighted) Nash welfare, if it achieves the highest Nash welfare among all allocations satisfying $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1$ for all $j\in{\mathcal{M}}$.
It is well known that the Nash welfare maximizing allocations correspond exactly to the optimal solutions of the convex program introduced by Eisenberg and Gale (1959).
$\max$ $\sum_{i\in{\mathcal{A}}}b_{i}\log u_{i}(\mathbf{y}_{i})\ \ \text{s.t.}$ $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1,\ \forall j\in{\mathcal{M}}\ \ \text{and}$ $\penalty 10000\ \penalty 10000\ \penalty 10000\ y_{ij}\geq 0,\ \forall i\in{\mathcal{A}},j\in{\mathcal{M}}\,.$ (EG)
@@
$\varphi(\mathbf{q})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{i\in{\mathcal{A}}}\max_{\mathbf{y}_{i}\in\mathbb{R}_{\geq 0}^{{\mathcal{M}}}}\left(b_{i}\log u_{i}(\mathbf{y}_{i})-\left\langle\mathbf{q},\mathbf{y}_{i}\right\rangle\right)+\sum_{j\in{\mathcal{M}}}q_{j}.$ (4)
We define the domain $\operatorname{dom}(\varphi)\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\cap_{i\in{\mathcal{A}}}\operatorname{pdom}(u_{i})$ (see (3)). Thus, $\operatorname{dom}(\varphi)$ is the set of nonnegative prices $\mathbf{q}$ such that $\varphi(\mathbf{q})<\infty$. Slater’s condition is satisfied; therefore strong duality holds, and the dual optimum is attained. Consequently, the optimal Nash welfare value is equal to
$\min_{\mathbf{q}\in\operatorname{dom}(\varphi)}\varphi(\mathbf{q})\,$, (5)
and we refer to the optimal $\mathbf{q}\in\operatorname{dom}(\varphi)$ as Gale prices. We note that these may not be unique. To justify this term, let us recall the definition of the Gale demand correspondence (2). The term corresponding to agent $i$ in the definition of $\varphi(\mathbf{q})$ is the value of the assignments in the Gale demand correspondence of this agent at prices $\mathbf{q}$ and budget $b_{i}$. We denote by $\mathcal{GD}_{i}(\mathbf{q},b_{i})$ the Gale demand corresponding to the utility function $u_{i}$, $i\in{\mathcal{A}}$.
@@ 2.4 Gale substitute utilities
    for any $b>0$, any $\mathbf{q},\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$, and any $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$, if $\mathbf{q}\geq\mathbf{q}^{\prime}$, then there exists a $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ such that $y^{\prime}_{j}\leq y_{j}$ whenever $q_{j}^{\prime}=q_{j}$;
@@
    for any $b\geq b^{\prime}>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$, and for any $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$, there exists a $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$ such that $\mathbf{y}^{\prime}\leq\mathbf{y}$.
We say that $u(\cdot)$ is $\Sigma$-Gale-substitutes, if it satisfies the Gale-substitutes condition along with the following additional property:
$\mbox{if }\mathbf{0}\in\operatorname{pdom}(u)\,,\,b>0\mbox{ and }\mathbf{q}\in\mathbb{R}_{\geq 0}^{\mathcal{M}},\mbox{ then for any }\mathbf{y}\in\mathcal{GD}(\mathbf{0},b),\exists\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b)\,:\,\mathbf{y}^{\prime}\leq\mathbf{y}\,.$ ($\Sigma$)
@@
As noted above, $\mathbf{0}\in\operatorname{pdom}(u)$ if and only if $u$ is satiable; thus, ($\Sigma$) is vacuously true for non-satiable utilities. In the satiable case, it asserts that if $\mathbf{y}$ is a bundle of maximum possible utility, then at any prices there is an optimal bundle below $\mathbf{y}$. Observe that this is immediate for separable utilities. On the other hand, ($\Sigma$) fails for the simple Leontief-free utility $u(\mathbf{x})=\max\{x_{1}+x_{2},1\}$: $\mathbf{y}=(1,0)$ is an optimal bundle at $\mathbf{q}=\mathbf{0}$, whereas for $b=1$, $\mathbf{q}=(2,1)$, the unique solution in $\mathcal{GD}^{u}(\mathbf{q},b)$ is $\mathbf{y}^{\prime}=(0,1)$.
Next, we present a utility function that satisfies the Gale-substitutes property but does not exhibit the weak gross substitutes (WGS) property.
@@
Consider an agent with budget $b$ and a utility function over two goods given by $u(\mathbf{x})=\min\{x_{1},0.4\}+x_{2}$. This utility is separable and, as such, Gale-substitutes (see Theorem 1.4). However, we show that this is not WGS: there exist prices $\mathbf{p}\leq\mathbf{p}^{\prime}$ and $\mathbf{x}\in{\mathcal{D}}^{u}(\mathbf{p},b)$ such that for some good $j$ with $p_{j}=p^{\prime}_{j}$, every $\mathbf{x}^{\prime}\in{\mathcal{D}}^{u}(\mathbf{p},b)$ has $x^{\prime}_{j}<x_{j}$. That is, let $\mathbf{p}=(0.8,1)$ and $\mathbf{p}^{\prime}=(0.9,1)$, and $b=1$. It is easy to see that $\mathbf{x}=(0.4,0.68)$ is the unique optimal bundle at prices $\mathbf{p}$, and $\mathbf{x}^{\prime}=(0.4,0.64)$ is the unique optimal bundle at prices $\mathbf{p}^{\prime}$ in contradiction to the WGS property.
We note that the gross substitutes property is one of the key conditions that allows for the efficient computation of market equilibria. However, SPLC utilities do not satisfy the gross substitutes property, as illustrated in the above example. In a later section, we show that SPLC preferences are Gale-substitutes, which forms the foundation of our main results.
3 The price of anarchy bound
We now show the ‘price of anarchy’ result stated in Theorem 1.1. The proof employs an argument based on concavity and the log-sum inequality, drawing inspiration from Theorem 1 in Brânzei et al. (2022).
Theorem 3.1 (Restatement of Theorem 1.1).
Given a Fisher market instance with concave utility functions, consider any competitive equilibrium $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$ and a Nash welfare maximizing allocation $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$. Then,
$\left(\prod_{i\in{\mathcal{A}}}u_{i}(\mathbf{x}_{i})^{b_{i}}\right)^{\frac{1}{\sum_{i\in{\mathcal{A}}}b_{i}}}\geq\left(\frac{1}{\mathrm{e}}\right)^{\frac{1}{\mathrm{e}}}\left(\prod_{i\in{\mathcal{A}}}u_{i}(\mathbf{y}_{i})^{b_{i}}\right)^{\frac{1}{\sum_{i\in{\mathcal{A}}}b_{i}}}.$
Proof.
Without loss of generality, we can assume $\sum_{i\in{\mathcal{A}}}b_{i}=1$. Note that since $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$ is a competitive equilibrium, then $u_{i}(\mathbf{x}_{i})\geq\min\left\{1,\frac{b_{i}}{\left\langle\mathbf{p},\mathbf{y}_{i}\right\rangle}\right\}u_{i}(\mathbf{y}_{i})$, as agent $i$ can use their budget, $b_{i}$, to buy at least $\frac{b_{i}}{\left\langle\mathbf{p},\mathbf{y}_{i}\right\rangle}$ fraction of $\mathbf{y}_{i}$. The lower bound holds because if $b_{i}<\left\langle\mathbf{p},\mathbf{y}_{i}\right\rangle$, then $u_{i}\left(\frac{b_{i}}{\left\langle\mathbf{p},\mathbf{y}_{i}\right\rangle}\mathbf{y}_{i}\right)\geq\frac{b_{i}}{\left\langle\mathbf{p},\mathbf{y}_{i}\right\rangle}u_{i}(\mathbf{y}_{i})$ using concavity and $u_{i}(\mathbf{0})=0$. Thus,
$\sum_{i\in{\mathcal{A}}}{b_{i}}\log\frac{u_{i}(\mathbf{y}_{i})}{u_{i}(\mathbf{x}_{i})}\leq\sum_{i\in{\mathcal{A}}}{b_{i}}\log\max\left\{1,\frac{\left\langle\mathbf{p},\mathbf{y}_{i}\right\rangle}{b_{i}}\right\}.$
Let $b^{\prime}_{i}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}{\left\langle\mathbf{p},\mathbf{y}_{i}\right\rangle}$, and let $\mathcal{A}$$^{\prime}$ be the set of agents such that $b^{\prime}_{i}\geq b_{i}$, $B\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{i\in\mathcal{A}^{\prime}}b_{i}$, and $B^{\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{i\in\mathcal{A}^{\prime}}b^{\prime}_{i}$. Therefore, using the log sum inequality,
$\sum_{i\in{\mathcal{A}}}{b_{i}}\log\frac{u_{i}(\mathbf{y}_{i})}{u_{i}(\mathbf{x}_{i})}\leq\sum_{i\in{\mathcal{A}}^{\prime}}{b_{i}}\log\frac{b^{\prime}_{i}}{b_{i}}\leq B\log\frac{B^{\prime}}{B}\,.$
By definition, $B\leq\sum_{i\in{\mathcal{A}}}b_{i}=1$. We also note that $B^{\prime}\leq\sum_{i\in{\mathcal{A}}}b_{i}^{\prime}=\sum_{i\in{\mathcal{A}}}\left\langle\mathbf{p},\mathbf{y}_{i}\right\rangle\leq\sum_{j\in{\mathcal{M}}}p_{j}\leq\sum_{i\in{\mathcal{A}}}b_{i}\leq 1$.
$\sum_{i\in{\mathcal{A}}}{b_{i}}\log\frac{u_{i}(\mathbf{y}_{i})}{u_{i}(\mathbf{x}_{i})}\leq\max_{B\leq 1}B\log\frac{1}{B}\leq\frac{1}{\mathrm{e}}\,.$
The statement follows. ∎
The following example demonstrates that the bound $(1/\mathrm{e})^{1/\mathrm{e}}$ is tight.
Example 3.2 (Tight example on price of anarchy).
Consider an instance with two agents and two goods. The first agent has a budget of $1$ and has a utility function:
$u_{1}(\mathbf{x}_{1})=x_{11}+\varepsilon x_{12},$
and the second agent has a budget of $\mathrm{e}-1$ and has a utility function
$u_{2}(\mathbf{x}_{2})=\varepsilon x_{21}+\min\{x_{22},1-\varepsilon\},$
where $\varepsilon>0$ is sufficiently small. The maximum Nash welfare is at least $(1-\varepsilon)^{(\mathrm{e}-1)/\mathrm{e}}$, taken at $\mathbf{y}_{1}=(1,0)$ and $\mathbf{y}_{2}=(0,1)$. Conversely, the unique competitive equilibrium is characterized by prices $p_{1}=\mathrm{e}(\frac{1}{1+\varepsilon})$ and $p_{2}=\mathrm{e}(\frac{\varepsilon}{1+\varepsilon})$, with corresponding allocations given by:
$\mathbf{x}_{1}=\left(\frac{1+\varepsilon-\mathrm{e}\varepsilon^{2}}{\mathrm{e}},\varepsilon\right),\quad\mathbf{x}_{2}=\left(1-\frac{1+\varepsilon-\mathrm{e}\varepsilon^{2}}{\mathrm{e}},1-\varepsilon\right).$
The Nash welfare objective for this allocation is $\left(\frac{1+\varepsilon}{\mathrm{e}}\cdot\left[1-\varepsilon+\varepsilon\cdot\left(1-\frac{1+\varepsilon-\mathrm{e}\varepsilon^{2}}{\mathrm{e}}\right)\right]^{\mathrm{e}-1}\right)^{\frac{1}{\mathrm{e}}}$. When $\varepsilon\to 0$, the Nash welfare ratio between these two allocations converges to $(1/\mathrm{e})^{1/\mathrm{e}}$.
The above example also shows that the price of stability is the same as the price of anarchy, as there is a unique CE.
4 Nash welfare approximates competitive equilibrium
We now derive Theorem 1.2, showing that Nash welfare maximizing allocations can be seen as approximate competitive equilibria. Our first simple lemma shows that the Gale demand solutions are feasible with respect to the budget: